# Judging It, Washing It: Scoring and Greenwashing Corporate Climate Disclosures using Large Language Models

### v1

@@ 1 Introduction
Corporate climate disclosures are a critical component of corporate climate action. They report information on their climate-related risks, emission reduction strategies and targets, and offer progress updates on a regular basis. Through these disclosures, corporations can provide transparency and accountability to their stakeholders, including investors, regulators, and consumers. Various reporting frameworks have been widely used, including CDP (formerly Carbon Disclosure Project), TCFD (Task Force on Climate-related Financial Disclosures), CSRD (Corporate Sustainability Reporting Directive), and efforts are underway to harmonize and standardize them. The number of reporting companies is growing rapidly. For example, the number of companies voluntarily disclosing to CDP increased from 2,600 in 2018 to 23,000 in 2023.
@@ 2 Related Work
There is a quickly growing body of literature on the use of Natural Language Processing (NLP) and machine learning methods to contribute to tackling climate change [2, 3]. They include efforts to detect, analyze, and fact-check environmental claims and stances [4, 5, 6, 7, 8, 9], identify topics and trends over time [10, 11], improve the performance of conversational AI agents with regards to climate change related information [12, 13, 14], and tools to support climate policymaking [15, 16].
There is also a number of recent works that employ LLMs to analyze environmental assessment reports, corporate sustainability reports, and corporate climate disclosure documents. The LLMs have proven themselves to be very versatile, capable of sifting through lengthy documents to detect specific items of interest, such as emission reduction targets [17] and sustainable development goals [18]. Furthermore, the LLMs can also be used to analyze entire reports to generate overall assessments of a company’s performance or transition plans [19, 20].
@@ 4 LLM-as-a-Judge (LLMJ) for Climate Disclosures
In both scoring types, we follow [14] in asking the LLM to consider accuracy, specificity, and completeness (“epistemological metrics”) and clarity (a “presentational metric”), in addition to the actual content of the response. We also specify factors that the LLM should not consider, such as the raw length of the response or irrelevant information. We use OpenAI’s GPT-4o-mini-2024-07-18 for our experiments, sampling with temperature parameter t=0.
@@ 4.2 Pairwise Comparison
In the pairwise comparison scoring system, we score a response by individually comparing it to $k$ other uniformly selected responses, asking the LLM to evaluate which response is “better” and ranking the response overall in terms of its “expected win rate,” out of 100%.
@@ 4.4 Evaluation
For the numerical rating system, we simply compute the weighted score (from 1-5) for all 1,416 companies. For the more costly pairwise comparison system, we evaluated all 147 A-List companies and randomly sampled 147 non-A-List companies, and performed $k=24$ comparisons against a random sample of the entire set.
@@ 4.5 Results
The two scoring systems produce consistent results. As shown in Figure 3, responses given a high rating by one scoring system are very likely to also have a high rating on the other.
Using at least one reference example is helpful. There is a clear increase in separation when going from zero-shot to one-shot prompting. However, going from one reference example to two does not clearly show any additional improvement. Using an indicative scale does not seem to improve separation, but does change the distribution of scores. Our particular choice of indicative scale shifted responses away from scores near 3 and towards scores near 2 and 4; this suggests that one can roughly tune the score distribution by carefully choosing the scale. For pairwise comparison scoring, chain-of-thought prompting is moderately helpful, but seems detrimental for numerical rating.
@@ 5.1 Greenwashing with LLMs
First, GPT-4o is, unsurprisingly, very good at greenwashing, particularly when it is allowed to hallucinate plans, goals, actions taken, and so on, generating responses that score an average of 0.63 points higher on the 5-point numerical rating scale and 14 points higher on the 100-point pairwise comparison scale. On the numerical rating scale, 55% of responses saw a score increase of at least half a point, and 17% saw a score increase of 1 point or more.
@@
    Adding completely false/hallucinated information, mostly about targets met (for example, “We reduced our Scope 2 emissions by 10% over the last year.”). We often observe the LLM changing planned actions (in the original response) to achieved actions (in the modified one).
@@ 6 Limitations and Future Work
Similarly, we ran our analysis on a single LLM (OpenAI’s ChatGPT-4o-mini). Given that LLMs continue to evolve and improve at a rapid pace, it would be valuable to repeat the analysis on other state-of-the-art LLMs and future generations of LLMs, so that we can gather more data points on the performance of the LLMJ method and its various in-context learning, indicative scale, and chain-of-thought techniques against different language models.
@@ 7 Conclusion
We focus on scoring disclosures on emission reduction targets and progress, which is perhaps the most tangible and direct way that a company’s climate action can be tracked and evaluated. Because the disclosures include both planned actions and actual progress, it is important to distinguish between them, and we find that the LLMJ methodology is capable of doing so.
@@ Appendix D Controlling for Length
To address the possibility of length being a confounding factor, we run a simple experiment in which we control for content while varying length. We use the same uniformly sampled set of 100 non-A-list companies from Section 5 and double the length of the companies’ responses by repeating the response twice. We then use the numerical rating system (with one-shot learning) to score these new responses. We compare to compare the original ratings to the new ratings in Figure 12. We see that nearly all scores are at or below the $y=x$ line, and in fact most of the points are below the line, indicating that doubling the length reduced the score.

### v2

@@ 1 Introduction
Corporate climate disclosures are a critical component of corporate climate action. They report information on their climate-related risks, emission reduction strategies and targets, and offer progress updates on a regular basis. Through these disclosures, corporations can provide transparency and accountability to their stakeholders, including investors, regulators, and consumers. Various reporting frameworks have been widely used, including CDP (formerly Carbon Disclosure Project), TCFD (Task Force on Climate-related Financial Disclosures), CSRD (Corporate Sustainability Reporting Directive), and efforts are underway to harmonize and standardize them. The number of reporting companies is growing rapidly. For example, the number of companies voluntarily disclosing to CDP increased from 2,600 in 2018 to 23,000 in 2023. The European Union is expecting 50,000 companies to report to CSRD in 2025.
@@ 2 Related Work
There is a quickly growing body of literature on the use of Natural Language Processing (NLP) and machine learning methods to contribute to tackling climate change [2, 3]. They include efforts to detect, analyze, and fact-check environmental claims and stances [4, 5, 6, 7, 8, 9, 10, 11], identify topics and trends over time [12, 13], improve the performance of conversational AI agents with regards to climate change related information [14, 15, 16], and tools to support climate policymaking [17, 18].
There is also a number of recent works that employ LLMs to analyze environmental assessment reports, corporate sustainability reports, and corporate climate disclosure documents. The LLMs have proven themselves to be very versatile, capable of sifting through lengthy documents to detect and extract specific items of interest, such as emission reduction targets [19, 20] and sustainable development goals [21]. Furthermore, the LLMs can also be used to analyze entire reports to generate overall assessments of a company’s performance or transition plans [22, 23].
@@ 4 LLM-as-a-Judge (LLMJ) for Climate Disclosures
In both scoring types, we follow [16] in asking the LLM to consider accuracy, specificity, and completeness (“epistemological metrics”) and clarity (a “presentational metric”), in addition to the actual content of the response. We also specify factors that the LLM should not consider, such as the raw length of the response or irrelevant information. [^1] We use OpenAI’s GPT-4o-mini-2024-07-18 for our experiments, sampling with temperature parameter t = 0. Our code is publicly available at https://github.com/mariannechuang/llm-corp-disclosure.
@@ 4.2 Pairwise Comparison
In the pairwise comparison scoring system, we score a response by individually comparing it to $k$ other uniformly selected responses, asking the LLM to evaluate which response is “better” and ranking the response overall in terms of its “expected win rate,” out of 100%. For example, a response which is rated “better” in 15 comparisons and “worse” in 5 would receive a score of 75 out of 100.
@@ 4.4 Evaluation
For the numerical rating system, we simply compute the weighted score (from 1-5) for all 1,416 companies. For the more costly pairwise comparison system, we evaluated all 147 A-List companies and randomly sampled 147 non-A-List companies, and performed $k=24$ comparisons against a random sample of the entire set. To make an apples-to-apples comparison, we bucket the numerical rating scores into 25 bins of equal width from 1 to 5.
@@ 4.5 Results
The two scoring systems produce consistent results. As shown in Figure 3, the scores given by the two systems are highly correlated, with $r^{2}=0.70$.
Using at least one reference example is helpful. There is a clear increase in separation when going from zero-shot to one-shot prompting. However, going from one reference example to two does not clearly show any additional improvement. Using an indicative scale does not seem to improve separation, but does change the distribution of scores. Our particular choice of indicative scale shifted responses away from scores near 3 and towards scores near 2 and 4 (see Figure 7(b) versus  7(e)); this suggests that one can roughly tune the score distribution by carefully choosing the scale. For pairwise comparison scoring, chain-of-thought prompting is moderately helpful, but seems detrimental for numerical rating.
@@ 5.1 Greenwashing with LLMs
First, GPT-4o-mini is, unsurprisingly, quite capable at greenwashing, particularly when it is allowed to hallucinate plans, goals, actions taken, and so on, generating responses that score an average of 0.63 points higher on the 5-point numerical rating scale and 14 points higher on the 100-point pairwise comparison scale. On the numerical rating scale, 55% of responses saw a score increase of at least half a point, and 17% saw a score increase of 1 point or more.
@@
    Adding completely false/hallucinated information, mostly about targets met (for example, “We reduced our Scope 2 emissions by 10% over the last year.”). This only occurs in the unconstrained case. In particular, we often observe the LLM changing planned actions (in the original response) to achieved actions (in the modified one).
@@ 6 Limitations and Future Work
Similarly, we ran our analysis on a single LLM (OpenAI’s ChatGPT-4o-mini). Given that LLMs continue to evolve and improve at a rapid pace, it would be valuable to repeat the analysis on other state-of-the-art LLMs and future generations of LLMs, so that we can gather more data points on the performance of the LLMJ method and its various in-context learning, indicative scale, and chain-of-thought techniques against different language models. By using different LLMs to evaluate LLM-generated greenwashed responses, one can also test for self-enhancement biases in the LLMJ methodology in this context.
@@ 7 Conclusion
We focus on scoring disclosures on emission reduction targets and progress, which is arguably the most tangible and direct way that a company’s climate action can be tracked and evaluated. Recognizing the fact that the claims made by the companies may not have been fact-checked by CDP, our analysis shows that the LLMJ methodology can effectively evaluate the claims when taken at face value.
@@
While the pairwise comparison scoring system outperformed numerical rating on the EMD metric, we must recognize that it incurs significantly higher computational costs (by a factor of $k$, the number of companies to compare against). Further, there is the practical issue of gaining access to responses from a representative set of companies, either from the current year, or from a previous reference year. This may be particularly challenging for individual companies, so an organization like CDP could consider sponsoring a benchmark dataset.
@@
Acknowledgments
G.C. acknowledges funding support by a U.S. National Science Foundation Graduate Research Fellowship.
@@ Appendix D Controlling for Length
To address the possibility of length being a confounding factor, we run a simple experiment in which we control for content while varying length. We use the same uniformly sampled set of 100 non-A-list companies from Section 5 and double the length of the companies’ responses by repeating the response twice. We then use the numerical rating system (with one-shot learning) to score these new responses. We compare the original ratings to the new ratings in Figure 12. We see that nearly all scores are at or below the $y=x$ line, and in fact most of the points are below the line, indicating that doubling the length reduced the score.